# On the topological ranks of Banach -algebras associated with groups of subexponential growth

### v2

@@ Abstract
[…]
@@ 1 Introduction
The (left) topological stable rank of a unital Banach algebra $\mathfrak{B}$ was introduced by Rieffel [Ri83] as the smallest $n\in\mathbb{Z}^{+}$ such that
@@
is dense in $\mathfrak{B}^{n}$. The idea was to produce a noncommutative notion of dimension similar to both the covering dimension for compact spaces and the Bass stable rank. This idea was then applied very successfully to the study of K-theory for $C^{*}$-algebras [Ri83b] and motivated the notion of real rank. Indeed, the real rank of a unital $C^{*}$-algebra $\mathfrak{B}$, as introduced by Brown and Pedersen [BP91], is the smallest $n\in\mathbb{N}$ such that
@@
Studying these noncommutative notions of dimension became a very interesting problem for $C^{*}$-algebraists and motivated many works within the area. This is particularly true in the case of group $C^{*}$-algebras and crossed products. Just to mention some examples, the topological stable rank has been studied for group $C^{*}$-algebras in [SuTa95, DyHaRo97], whereas the real rank of group $C^{*}$-algebras has been studied in [Ka93, ArKa12, Mo23]. In the setting of crossed products, we can also find the works [Pu90b, Pu90, CP05].
There are also many works on topological ranks that stay outside the realm of $C^{*}$-algebras (see [Ba98, Ba99, DaJi09]) and this will be our case too. Indeed, we will be interested in the topological ranks of (generalized) convolution algebras and $L^{p}$-operator algebras, as introduced by Christopher Phillips [CP12, CP13].
The study of $L^{p}$-operator algebras has attracted much attention in the last decade (see [Ga21] and the references therein) and it is still a very active area of research [BiDeWe24, ChGaTh24, el24]. Most of the articles seem to be concerned with the properties that work exactly like in the $C^{*}$-case [el24] and a few others deal with new (perhaps surprising) phenomena [BiDeWe24, ChGaTh24]. In this sense, the present article should be put in the first group, as we plan to show that the topological stable rank is independent of $p$, for a large class of symmetrized crossed products. Symmetrized $L^{p}$-crossed products are ‘-analogs’ of regular $L^{p}$-crossed products and have the advantage of being Banach -algebras naturally acting on a natural $L^{2}$-space. They have been studied in [CP19, SaWi20, AuOr22, el24b, el24, BaKw25], often under the name of ‘algebras of pseudofunctions’.
In the case of topological stable rank 1, we can also say something about regular $L^{p}$-crossed products. In fact, we also aim to solve the following question of Christopher Phillips [CP13, Question 8.4].
@@
Instead of $C^{*}$-dynamical systems, we chose to work with Fell bundles over groups of subexponential growth. This is in order to maximize generality, but also because the main tools we used were already developed in this framework [BeCo24, Fl24]. Note that the inclusion of Fell bundles provides almost no extra complication (when the approach is compared to, say, an approach based strictly on twisted $C^{*}$-dynamical systems). We also hope that this generality will be useful in the future.
[…]
@@
[…]
@@
[…]
The class of Banach -algebras sitting between $\ell^{1}(G\,|\,{\mathscr{C}})$ and ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ seems to be vast. Indeed, for twisted $C^{*}$-dynamical systems of the form $(G,C_{0}(X),\alpha,\omega)$ and taking a subset $P\subset[1,\infty]$ as an input, Bardadyn and Kwaśniewski constructed an ‘$L^{P}$-completion’ of $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$. If $P$ is symmetric with respect to $2$, this is a Banach -algebra [BaKw25, Remark 4.13], which embeds in $C_{0}(X)\rtimes_{\alpha,\omega}G$ by the Riesz-Thorin interpolation theorem.
@@
In fact, and thanks to the generality of our setting, a lot of other topological ranks are computed in Theorem 3.2. They might be regarded as $L^{p}$-versions of results in [Pu90b, Pu90, BlKuRo92, Ka93, CP05]. It was recently announced by Li and Niu that every crossed product $C(X)\rtimes\mathbb{Z}^{d}$, coming from a free and minimal action $\mathbb{Z}^{d}\curvearrowright X$, has topological stable rank $1$ [LiNi23]. This widens the class of algebras where our theorems are applicable and provides other examples not explicitly mentioned here.
The reader should notice that some of the methods used in these papers are very different and some results are very intricate, but we succeeded in bypassing these methods.
The paper is organized as follows. Section 2 deals with preliminaries but also the proof of Theorem 1.2. It first introduces Fell bundles and their $\ell^{1}$-algebras and recalls some of the tools that will be later involved in the proofs of the main results. This section is then divided into two subsections. The first subsection introduces and studies the topological stable rank for these Fell bundle algebras, while the second subsection does the same for the real rank. Section 3 then briefly discusses the construction of $L^{p}$-crossed products and symmetrized $L^{p}$-crossed products and explains how the previously obtained results apply to these algebras. The main theorem of this section is Theorem 3.2, which basically consists of a list of results that follow from Theorem 3 and their (known) appropriate $C^{*}$-counterparts.
@@ 2 $\ell^{1}$-algebras of Fell bundles
From now on, $G$ will denote an amenable group with unit ${\sf e}$. We will also fix a Fell bundle $\mathscr{C}\!=\bigsqcup_{x\in G}\mathfrak{C}_{x}$ over $G$. The associated algebra of summable cross-sections $\ell^{1}(G\,|\,\mathscr{C})$ is a Banach -algebra and a completion of the space $C_{\rm c}(G\,|\,\mathscr{C})$ of cross-sections with finite support under the norm
@@
The universal $C^{*}$-algebra completion of $\ell^{1}(G\,|\,\mathscr{C})$ is denoted by ${\rm C^{*}}(G\,|\,\mathscr{C})$ and the associated universal $C^{*}$-norm is denoted $\lVert\cdot\rVert_{*}$. For the general theory of Fell bundles, we recommend [FD88, Chapter VIII]. Here, we only recall the product on $\ell^{1}(G\,|\,\mathscr{C})$, given by
@@
primarily because $\lVert\Phi\rVert_{\infty}\leq\lVert\Phi\rVert_{*}$, for all $\Phi\in{\rm C^{*}}(G\,|\,\mathscr{C})$ [Fl24, Lemma 3.4].
@@
Let $\mathfrak{A}$ be a $C^{*}$-algebra. A (continuous) twisted action of $G$ on $\mathfrak{A}$ is a pair $(\alpha,\omega)$ of maps $\alpha:G\to{\rm Aut}({\mathfrak{A}})$, $\omega:G\times G\to\mathcal{UM}({\mathfrak{A}})$, such that
@@
making $\ell^{1}_{\alpha,\omega}(G,\mathfrak{A})$ a Banach -algebra under the norm $\lVert\Phi\rVert_{1}=\sum_{x\in G}\lVert\Phi(x)\rVert_{\mathfrak{A}}$. When the twist is trivial ($\omega\equiv 1$), we omit any mention of it and call the resulting algebra $\ell^{1}_{\alpha}(G,\mathfrak{A})$ as the convolution algebra associated with the action $\alpha$. In this case, the triple $(G,\mathfrak{A},\alpha)$ is called a (untwisted) $C^{*}$-dynamical system. The associated $C^{*}$-envelope is usually denoted $\mathfrak{A}\rtimes_{\alpha,\omega}G$, or $\mathfrak{A}\rtimes_{\alpha}G$ in the untwisted case.
It is well known that these algebras correspond to our algebras $\ell^{1}(G\,|\,\mathscr{C})$, for very particular Fell bundles. In fact, these bundles may be easily described as $\mathscr{C}=\mathfrak{A}\times G$, with quotient map $q(a,x)=x$, constant norms $\lVert\cdot\rVert_{\mathfrak{C}_{x}}=\lVert\cdot\rVert_{\mathfrak{A}}$, and operations
@@
In what follows, $\widetilde{\mathfrak{B}}$ will denote the minimal unitization of $\mathfrak{B}$ (if $\mathfrak{B}$ was already unital, we simply set $\widetilde{\mathfrak{B}}=\mathfrak{B}$). We note that the minimal unitization of $\ell^{1}(G\,|\,\mathscr{C})$ corresponds to the $\ell^{1}$-algebra of the ‘unitization Fell bundle’ $\widetilde{\mathscr{C}}\to G$, which is constructed from $\mathscr{C}$ simply by replacing $\mathfrak{C}_{\sf e}$ with $\widetilde{\mathfrak{C}_{\sf e}}$ and keeping all the other fibers intact.
In order to prove our main result, we first need to recall some notions of multipliers. In particular, the following is a well-known characterization of amenability that can be found in [BrOz08, Theorem 2.6.8]. It uses the notion of positive definite functions, so we recall that first.
@@
A function $\varphi:G\to\mathbb{C}$ is said to be positive definite if the matrix
$[\varphi(x^{-1}y)]_{x,y\in K}\in\mathbb{M}_{|K|}(\mathbb{C})$
is positive definite for every finite subset $K\subset G$.
@@
$G$ is amenable if and only if there exists a net $\{\varphi_{j}\}_{j\in J}$ of finitely supported positive definite functions on $G$ such that $\varphi_{j}\to 1$ pointwise and $\varphi_{i}({\sf e})\leq 1$, for all $j\in J$.
Furthermore, Bédos and Conti [BeCo24, Example 4.5] showed that a positive definite function $\varphi:G\to\mathbb{C}$ induces a c.p. map
@@
In fact, it is easy to see that, if $\varphi_{j}\to 1$ pointwise and $\varphi_{j}({\sf e})\leq 1$ (as in the previous lemma), then $\lVert T_{\varphi_{j}}\rVert\leq 1$ and $T_{\varphi_{j}}\Phi\to\Phi$, for all $\Phi\in C_{\rm c}(G\,|\,\mathscr{C})$.
The other ingredient we need to introduce comes from spectral theory. So, as usual, we will denote the spectrum of an element $b\in\mathfrak{B}$ by
${\rm Spec}_{\mathfrak{B}}(b)=\{\lambda\in\mathbb{C}\mid b-\lambda 1\textup{ isnot invertible in }\widetilde{\mathfrak{B}}\}.$
We also recall that a group is said to have subexponential growth if
$\lim_{n\to\infty}|K^{n}|^{1/n}=1,$
for all finite subsets $K\subset G$. All nilpotent groups have subexponential growth, so they serve as examples. In fact, all of our concrete applications are about nilpotent groups (see Section 3).
Theorem 2.4 ([Fl24]).
@@
Before moving to the next section, we remark that all the Banach -algebras mentioned in this section have a continuous (actually isometric) involution. In what follows, the algebras that admit a continuous involution will play a central role.

### v3

@@ Abstract
Let $G$ be a group of subexponential growth and $\mathscr{C}\overset{q}{\to}G$ a Fell bundle. We show that any Banach -algebra that sits between the associated $\ell^{1}$-algebra $\ell^{1}(G\,|\,\mathscr{C})$ and its $C^{*}$-envelope has the same topological stable rank and real rank as $\ell^{1}(G\,|\,\mathscr{C})$. We apply this result to compute the topological stable rank and real rank of various classes of symmetrized twisted $L^{p}$-crossed products and show that some twisted $L^{p}$-crossed products have topological stable rank 1. Our results are new even in the case of (untwisted) group algebras.
@@ 1 Introduction
The (left) topological stable rank of a unital Banach algebra $\mathfrak{B}$ was introduced by Rieffel [33] as the smallest $n\in\mathbb{Z}^{+}$ such that
@@
is dense in $\mathfrak{B}^{n}$. The idea was to produce a noncommutative notion of dimension similar to both the covering dimension for compact spaces and the Bass stable rank. This idea was then applied very successfully to the study of K-theory for $C^{*}$-algebras [34] and motivated the notion of real rank. Indeed, the real rank of a unital $C^{*}$-algebra $\mathfrak{B}$, as introduced by Brown and Pedersen [9], is the smallest $n\in\mathbb{N}$ such that
@@
Studying these noncommutative notions of dimension has become a very interesting problem for $C^{*}$-algebraists and has motivated many works in this area. This is particularly true in the case of group $C^{*}$-algebras and crossed products. Just to mention some examples, the topological stable rank has been studied for group $C^{*}$-algebras in [37, 14, 20, 32], whereas the real rank of group $C^{*}$-algebras has been studied in [22, 1, 36, 25]. In the setting of crossed products, we can also find the works [30, 31, 26].
There are also many works on topological ranks that remain outside the realm of $C^{*}$-algebras (see [3, 4, 12]), and this will be our case as well. Indeed, we will be interested in the topological ranks of generalized convolution algebras and $L^{p}$-operator algebras, as introduced by N. C. Phillips [27, 28].
The study of $L^{p}$-operator algebras has attracted much attention in the last decade (see [19] and the references therein), and it is currently a very active area of research [8, 11, 16]. Most of the articles seem to be concerned with the properties that work exactly as in the $C^{*}$-case [16], while a few others deal with new (perhaps surprising) phenomena [8, 11]. In this sense, the present article should be placed in the first group, as we plan to demonstrate that the topological stable rank is independent of $p$ for a large class of symmetrized crossed products. Symmetrized $L^{p}$-crossed products are ‘-analogs’ of regular $L^{p}$-crossed products and have the advantage of being Banach -algebras that naturally act on a $L^{2}$-space. They have been studied in [29, 35, 2, 15, 16, 5], often under the name of ‘algebras of pseudofunctions’.
In the case of topological stable rank 1, we can also say something about regular $L^{p}$-crossed products. In fact, we also aim to solve the following question of N. C. Phillips [28, Question 8.4].
@@
Instead of $C^{*}$-dynamical systems, we chose to work with Fell bundles over groups of subexponential growth. This is in order to maximize generality, but also because the main tools we used were already developed in this framework [6, 18]. Note that the inclusion of Fell bundles provides almost no extra complication (when the approach is compared to, say, an approach based strictly on twisted $C^{*}$-dynamical systems). We also hope that this generality will be useful in the future.
Now let us present the main theorem. Given a Fell bundle $\mathscr{C}\overset{q}{\to}G$, we let $\ell^{1}(G\,|\,{\mathscr{C}})$ be the associated algebra of summable cross-sections, and we write ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ for the associated enveloping $C^{*}$-algebra. Naturally, $\ell^{1}(G\,|\,{\mathscr{C}})$ is equipped with its $1$-norm, whereas ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ is considered with its $C^{*}$-norm. In what follows, we use the symbols ${\rm tsr}(\mathfrak{B})$ and ${\rm rr}(\mathfrak{B})$ to denote the topological stable rank and the real rank of the algebra $\mathfrak{B}$.
@@
Let $G$ be a group of subexponential growth and $\mathscr{C}\overset{q}{\to}G$ a Fell bundle. Suppose that $\mathfrak{B}$ is a Banach -algebra with a continuous involution and such that $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}\subset{\rm C^{*}}(G\,|\,{\mathscr{C}})$. Then
@@
Let $G$ be a group of subexponential growth and $\mathscr{C}\overset{q}{\to}G$ a Fell bundle. Let $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}$ be any continuous, dense inclusion into a Banach algebra $\mathfrak{B}$. Then
${\rm tsr}(\mathfrak{B})\leq{\rm tsr}\big({\rm C^{*}}(G\,|\,{\mathscr{C}})\big).$
In particular, if ${\rm tsr}\big({\rm C^{*}}(G\,|\,{\mathscr{C}})\big)=1$, then ${\rm tsr}(\mathfrak{B})=1.$
The class of Banach -algebras sitting between $\ell^{1}(G\,|\,{\mathscr{C}})$ and ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ seems to be vast. Indeed, for twisted $C^{*}$-dynamical systems of the form $(G,C_{0}(X),\alpha,\omega)$, and taking a subset $P\subset[1,\infty]$ as input, Bardadyn and Kwaśniewski constructed an $L^{P}$-completion’ of $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$. If $P$ is symmetric with respect to $2$, this is a Banach -algebra [5, Remark 4.13], which embeds in $C_{0}(X)\rtimes_{\alpha,\omega}G$ by the Riesz-Thorin interpolation theorem.
@@
In fact, thanks to the generality of our setting, many other topological ranks are computed in Theorem 3.3. They might be regarded as $L^{p}$-versions of results in [30, 31, 7, 22, 26]. It was recently announced by Li and Niu that every crossed product $C(X)\rtimes\mathbb{Z}^{d}$, coming from a free and minimal action $\mathbb{Z}^{d}\curvearrowright X$, has topological stable rank $1$ [23]. This widens the class of algebras for which our theorems are applicable and provides other examples not explicitly mentioned here.
We remark that the results mentioned in the previous paragraph involve many different approaches, some of which are very intricate, while others contain very ad-hoc techniques. Our approach is very general and manages to bypass these methods in a unified manner.
The reader should note that groups of subexponential growth are necessarily amenable; hence, we do not know what the situation is for non-amenable groups. More concretely, we do not know the topological stable rank or the real rank of $F^{p}_{*}(G)$ for non-amenable $G$. This question seems particularly interesting for free groups $G=\mathbb{F}_{n}$.
The article is organized as follows. Section 2 deals with preliminaries but also includes the proof of Theorem 1.2. It first introduces Fell bundles and their $\ell^{1}$-algebras and recalls some of the tools that will be later involved in the proofs of the main results. This section is divided into two subsections. The first subsection introduces and studies the topological stable rank for these Fell bundle algebras, while the second subsection does the same for the real rank. Section 3 then briefly discusses the construction of $L^{p}$-crossed products and symmetrized $L^{p}$-crossed products and explains how the previously obtained results apply to these algebras. The main theorem of this section is Theorem 3.3, which basically consists of a list of results that follow from Theorem 1.4 and their (already known) appropriate $C^{*}$-counterparts.
@@ 2 $\ell^{1}$-algebras of Fell bundles
From now on, $G$ will denote a (discrete) amenable group with unit $\sf e$}. We will also fix a Fell bundle $\mathscr{C}\!=\bigsqcup_{x\in G}\mathfrak{C}_{x}$ over $G$. That is, the $\mathfrak{C}_{x}$’s are a family of Banach spaces equipped with bilinear, associative multiplication maps $\bullet:\mathfrak{C}_{x}\times\mathfrak{C}_{y}\to\mathfrak{C}_{xy}$ and conjugate-linear, anti-multiplicative involutions $^{\bullet}:\mathfrak{C}_{x}\to\mathfrak{C}_{x^{-1}}$ satisfying natural properties that turn the unit fiber $\mathfrak{C}_{\sf e}$ into a $C^{*}$-algebra, and each $\mathfrak{C}_{x}$ into a Hilbert $\mathfrak{C}_{\sf e}$-bimodule (see [17]).
The associated algebra of summable cross-sections $\ell^{1}(G\,|\,\mathscr{C})$ is a Banach -algebra and a completion of the space $C_{\rm c}(G\,|\,\mathscr{C})$ of cross-sections with finite support under the norm
@@
The universal $C^{*}$-algebra completion of $\ell^{1}(G\,|\,\mathscr{C})$ is denoted by ${\rm C^{*}}(G\,|\,\mathscr{C})$, and the associated universal $C^{*}$-norm (i.e., the norm in ${\rm C^{*}}(G\,|\,\mathscr{C})$) is denoted $\lVert\cdot\rVert_{*}$. We recall that the product on $\ell^{1}(G\,|\,\mathscr{C})$ is given by
@@
primarily because $\lVert\Phi\rVert_{\infty}\leq\lVert\Phi\rVert_{*}$, for all $\Phi\in{\rm C^{*}}(G\,|\,\mathscr{C})$ [18, Lemma 3.4].
@@
Let $\mathfrak{A}$ be a $C^{*}$-algebra. A (continuous) twisted action of $G$ on $\mathfrak{A}$ is a pair $(\alpha,\omega)$ of maps $\alpha:G\to{\rm Aut}({\mathfrak{A}})$ and $\omega:G\times G\to\mathcal{UM}({\mathfrak{A}})$, such that $\alpha$ is strongly continuous, $\omega$ is jointly strictly continuous, and
@@
making $\ell^{1}_{\alpha,\omega}(G,\mathfrak{A})$ a Banach -algebra under the norm $\lVert\Phi\rVert_{1}=\sum_{x\in G}\lVert\Phi(x)\rVert_{\mathfrak{A}}$. When the twist is trivial ($\omega\equiv 1$), we omit any mention of it and call the resulting algebra $\ell^{1}_{\alpha}(G,\mathfrak{A})$ the convolution algebra associated with the action $\alpha$. In this case, the triple $(G,\mathfrak{A},\alpha)$ is called an (untwisted) $C^{*}$-dynamical system. The associated $C^{*}$-envelope is usually denoted $\mathfrak{A}\rtimes_{\alpha,\omega}G$ or $\mathfrak{A}\rtimes_{\alpha}G$ in the untwisted case.
It is well known that these algebras correspond to our algebras $\ell^{1}(G\,|\,\mathscr{C})$ for very particular Fell bundles. In fact, these bundles may be easily described as $\mathscr{C}=\mathfrak{A}\times G$, with the quotient map $q(a,x)=x$, constant norms $\lVert\cdot\rVert_{\mathfrak{C}_{x}}=\lVert\cdot\rVert_{\mathfrak{A}}$, and operations
@@
In what follows, $\widetilde{\mathfrak{B}}$ will denote the minimal unitization of $\mathfrak{B}$. If $\mathfrak{B}$ is already unital, we simply set $\widetilde{\mathfrak{B}}=\mathfrak{B}$. We note that the minimal unitization of $\ell^{1}(G\,|\,\mathscr{C})$ corresponds to the $\ell^{1}$-algebra of the ‘unitization Fell bundle’ $\widetilde{\mathscr{C}}\to G$, which is constructed from $\mathscr{C}$ simply by replacing $\mathfrak{C}_{\sf e}$ with $\widetilde{\mathfrak{C}_{\sf e}}$ and keeping all the other fibers intact.
In order to prove our main result, we first need to recall some notions of multipliers. In particular, the following is a well-known characterization of amenability that can be found in [10, Theorem 2.6.8]. It uses the notion of positive definite functions, so we will recall that first.
@@
Let $G$ be a group. A function $\varphi:G\to\mathbb{C}$ is said to be positive definite if, for every finite subset $K\subset G$, the matrix
$\varphi(x^{-1}y)]_{x,y\in K}\in\mathbb{M}_{|K|}(\mathbb{C})$
is positive definite.
@@
A group $G$ is amenable if and only if there exists a net $\{\varphi_{j}\}_{j\in J}$ of finitely supported positive definite functions on $G$ such that $\varphi_{j}\to 1$ pointwise and $\varphi_{j}({\sf e})\leq 1$, for all $j\in J$.
Furthermore, Bédos and Conti [6, Example 4.4] showed that a positive definite function $\varphi:G\to\mathbb{C}$ induces a completely positive map
@@
In fact, if $\varphi_{j}\to 1$ pointwise and $\varphi_{j}({\sf e})\leq 1$ (as in the previous lemma), then $\sup_{x\in G}|\varphi(x)|\leq 1$ [6, Corollary 3.7]. Moreover, one has $\lVert T_{\varphi_{j}}\Phi\rVert_{*}\leq\lVert\Phi\rVert_{*}$ and $T_{\varphi_{j}}\Phi\to\Phi$, for all $\Phi\in C_{\rm c}(G\,|\,\mathscr{C})$ [6, Theorem 3.11].
The other ingredient we need to introduce comes from spectral theory. Hence, as usual, we will denote the spectrum of an element $b\in\mathfrak{B}$ by
${\rm Spec}_{\mathfrak{B}}(b)=\{t\in\mathbb{C}\mid b-t1\textup{ is not invertible in }\widetilde{\mathfrak{B}}\}$.
We also recall that a group is said to have subexponential growth if, for all finite subsets $K\subset G$, one has
$\lim_{n\to\infty}|K^{n}|^{1/n}=1.$
All nilpotent groups have subexponential growth; thus, they serve as examples. In fact, all of our concrete applications make use of nilpotent groups (see Section 3).
Theorem 2.4 ([18]).
@@
Before moving to the next section, we remark that all the Banach -algebras mentioned in this section have a continuous (actually, isometric) involution. In what follows, the algebras that admit a continuous involution will play a central role.